OpenStudy (anonymous):

Solve the equation. y + (–6) = –9 A. –15 B. –3 C. 54 D. 3

OpenStudy (anonymous):

b :)..

OpenStudy (anonymous):

y+(-6)=-9. is y-6=-9. add 6 to both sides of the equation giving: y-6(+6)=-9+6

OpenStudy (anonymous):

did u want it explained? or just the answer?

OpenStudy (anonymous):

which is: y=-3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

your welcome!